# tenSVD algorithm for compression

### v1

@@ Abstract
Tensors provide a robust framework for managing high-dimensional data. Consequently, tensor analysis has emerged as an active research area in various domains, including machine learning, signal processing, computer vision, graph analysis, and data mining. This study introduces an efficient image storage approach utilizing tensors, aiming to minimize memory to store, bandwidth to transmit and energy to processing. The proposed method organizes original data into a higher-order tensor and applies the Tucker model for compression. Implemented in R, this method is compared to a baseline algorithm. The evaluation focuses on efficient of algorithm measured in term of computational time and the quality of information preserved, using both simulated and real datasets. A detailed analysis of the results is conducted, employing established quantitative metrics, with significant attention paid to sustainability in terms of energy consumption across algorithms.
keywords:
Tucker decomposition , Tensor compression , Tucker3 , HOSVD , Image and video processing , Sustainability
journal: Knowledge and Information Systems
@@ 1 Introduction
A significant volume and complexity of image and video datasets are continuously generated by modern devices every second. These data necessitate continuous improvement of methods to reduce storage space and processing time, thereby facilitating more efficient extraction and transmission of information. It is well known that there are two major categories of methods to reduce memory space, with compression being one. The first category is based on a lossless compression strategy, while the second relies on lossy compression. The latter guarantees substantial reductions in memory cost while maintaining a high percentage of the original information. Ongoing research endeavors persist in exploring and developing innovative methods to address the challenges associated with compression tasks.
@@
To this end, a diverse array of methodologies has been proposed across various disciplines, including astronomy, biology, chemistry, engineering, geology, and medical imaging. Several comprehensive surveys have been conducted to organize the different approaches employed in these domains, notably [2] and [3], which are recent surveys comparing several techniques for image and video compression.
@@
Image compression is effectively achieved using multivariate techniques such as Singular Value Decomposition (SVD) [4], [5]. While grayscale images are typically represented as two-dimensional matrices, color images are commonly structured in three dimensions (RGB channels). These multivariate methods can be applied to diminish the dimensionality of each color channel independently. However, such approaches become inefficient when dealing with videos or more complex signals that necessitate consideration of higher-dimensional data.
Consequently, methodologies such as Tucker analysis [6] and broader tensor decompositions are increasingly leveraged for compression purposes. For an overview of tensor decomposition models, refer to [7]. These approaches are better equipped to manage the complexity inherent in high-quality images or videos and their conversion into higher dimensional data, thereby enhancing both compression efficiency and overall performance.
@@
A significant bottleneck for all compression methods is achieving the extraction of only the useful information contained in high-resolution images or videos. In this way, a low storage footprint is needed while maintaining an acceptable level of quality. The trade-off between reduced storage and the details and clarity of images or videos is consistently considered in compression contexts. Additionally, the time required for compression processes is a critical factor in signal transmission, as it directly influences transmission efficiency and bandwidth utilization. Rarely is compression studied as a fundamental tool for achieving a more sustainable processing of images and videos. However, concerns regarding the energy consumption of data centers are rapidly increasing, as much of the energy they consume is for processing and archiving the millions of images and videos collected daily. Thus, more attention needs to be paid when developing algorithms for image and video compression.
@@
Based on tensor decomposition, a novel algorithm is proposed for this purpose, called tenSVD (tensor Singular Value Decomposition). It is designed to ensure high sustainability in terms of energy consumption while balancing the trade-off between storage and quality requirements.
@@
To determine effectiveness, the tenSVD is compared to the established algorithm for Tucker analysis, High Order Singular Value Decomposition (HOSVD) [8]. The strengths and weaknesses of tenSVD are analyzed using several metrics, including Peak Signal-to-Noise Ratio (PSNR), Mean Squared Error (MSE), and Processing Time (TIME).
@@
The remainder of this article is structured as follows: Section 2 introduces tensors, including preliminary concepts. Section 3 presents the main idea of the proposed approach and the relative algorithm. A case study with full comparative results is provided in Section 4. Finally, conclusions and findings are presented in Section 5.
@@ 2 Tensor Decomposition
The Tucker model (Tucker3) is one of the most cited and widely applied methods for tensor decomposition. Originally proposed for three-way cases, it is referred to in several ways, including Three-Mode Principal Component Analysis (3MPCA), N-Mode Components Analysis (N-Mode PCA), and Higher-Order SVD (HOSVD) [6, 9, 10, 8]. Today, the Tucker model has been generalized to higher-order tensors and is rapidly increasing its application in imaging analysis [11]. For clarity, some preliminary concepts and notations will be recalled in Section 2.1, which will be used throughout the paper.
@@ 2.1 Preliminary Concepts and Notations
Objects with more than two indices are referred to as multi-way arrays, whereby the number of indices is termed the order. Nowadays, it is common to refer to any multi-way array as a tensor, and we maintain this tradition to identify the linear operators along with their multidimensional counterparts [12].
A real-valued tensor $\mathcal{X}\in\mathbb{R}$ of order $N\in\mathbb{N}$ is denoted by indices $(I_{1}\times\dots\times I_{N})$, and its entries are denoted by $x_{i_{1}\dots i_{N}}$. When only a fixed subset of indices is considered, this part of the tensor is termed a subtensor. The fibers are defined by fixing all indices except for one, e.g., $\mathbf{x}(:,i_{2},i_{3},\dots,i_{N})$, and slices are obtained by fixing all but two indices, e.g., $\mathbf{X}(:,:,i_{3},\dots,i_{N})$.
Tensors can be reformatted in several ways, with one particular case of manipulation known as n-mode unfolding or matrix unfolding. Denoted by $\mathbf{X}{(n)}$, the n-mode unfolding of a tensor $\mathcal{X}$ arranges the n-mode fibers as columns in the resulting matrix: $\mathbf{X}{(n)}\ (I_{n}\times\prod_{i\neq n}I_{i})$.
According to Kolda and Bader [13], tensor products involved are significantly more complex compared to matrices.
Let $\mathcal{X}(I_{1}\times\dots\times I_{n}\times\dots\times I_{N})$ and $\mathbf{U}(H\times I_{n})$ be defined as follows:
@@
where $\mathcal{Y}(I_{1}\times\dots\times H\times\dots\times I_{N})$, and ×$\times_{n}$ is the n-mode product of a tensor.
@@
The inner product of two same-sized tensors $\mathcal{X}$ and $\mathcal{Y}$ is the sum of the products of their corresponding entries:
$\left\langle\mathcal{X},\ \mathcal{Y}\right\rangle=\sum_{i_{1}=1}^{I_{1}}\sum_{i_{2}=1}^{I_{2}}\dots\sum_{i_{N}=1}^{I_{N}}x_{i_{1}i_{2}\dots i_{N}}y_{i_{1}i_{2}\dots i_{N}}$
@@ 2.2 TUCKER MODEL
In accordance with the notation given in Section 2.1, the Tucker model for an Nth order tensor can be expressed as
@@
where $(n=1,2,\dots,N)$ and each $\mathbf{U}^{(n)}(I_{n}\times R_{n})$ is an orthonormal component matrix $(\mathbf{U}^{(n)t}\mathbf{U}^{(n)}=\mathbf{I})$; $\mathcal{G}$ is the core tensor with dimension $(R_{1}\times R_{2}\times\dots\times R_{N})$.
Assuming $N=2$ with dimension $(I_{1}\times I_{2})$, and $R_{1}$ and $R_{2}$ representing the dimensions of the vector space spanned by the first and second modes, respectively, the Tucker decomposition can be written as
@@
where $\mathbf{U}^{(1)}(I_{1}\times R_{1})$ and $\mathbf{U}^{(2)}(I_{2}\times R_{2})$ are orthonormal component matrices, and $\mathcal{G}$ is the core tensor with dimension $(R_{1}\times R_{2})$.
It can be demonstrated that $\mathbf{U}^{(1)}$ and $\mathbf{U}^{(2)}$ are the left and right singular vectors of $\mathcal{X}$, while the core tensor is a matrix that contains singular values on the first r diagonal elements ($r=\min(R_{1},R_{2}))$ and zeros otherwise.
It is important to underline that the core tensor for $N>2$ is a full tensor, where its elements can also be negative. The magnitude of these elements indicates the strength of the interaction between components represented by the corresponding indices in the component matrices.
@@ 2.3 Tucker Parameters Estimation
The Tucker decomposition of a 2nd order tensor is equivalent to a Singular Value Decomposition (SVD). Thus, it is not surprising that Higher-Order Singular Value Decomposition (HOSVD) is the most popular algorithm for estimating the parameters of a Tucker model in cases where $N>2$.
Introduced by De Lathauwer et al. [8], which generalizes the three-way case presented by Tucker [6], HOSVD finds the exact Tucker decomposition of rank-$(R_{1},R_{2},\dots,R_{N})$ and subsequently provides the core tensor $\mathcal{G}$. In essence, HOSVD initially computes the left singular vectors of $\mathbf{X}_{(n)}$ for $n=1,2,\dots,N$. Finally, the transposes of these matrices are multiplied with the original tensor $\mathcal{X}$ to obtain the core tensor.
When exact parameters are not required, such as in the case of lossy compression, only the first $r_{n}$ columns of $\mathbf{U}^{(n)}$ are considered. The objective function to solve is defined as follows:
@@
where $\mathbf{U}^{(n)}(I_{n}\times r_{n})$ is orthonormal, and $\mathcal{G}$ has dimensions $(r_{1}\times r_{2}\times\dots\times r_{N})$.
The details of the optimization problem are readily available in [8, 13]. The version of HOSVD used to estimate the parameters is referred to as truncated HOSVD (t-HOSVD), which is briefly described below, highlighting its simplicity and limited number of steps.
@@
Based on a lossless compression strategy, HOSVD can significantly reduce the memory cost of the original tensor. Its effectiveness is determined by $\text{rank}-(R_{1},R_{2},\dots,R_{N})$ compared to the dimensions of the modes $(I_{1},I_{2},\dots,I_{N})$. In fact, the original memory cost of $\mathcal{X}$ is $\prod_{n=1}^{N}I_{n}$, while the memory cost of HOSVD is given by the sum of smaller matrices plus a core tensor, specifically $\sum_{n=1}^{N}I_{n}\times R_{n}$ for the component matrices and $\prod_{n=1}^{N}R_{n}$ elements for core tensor.
Differently, by choosing $r_{1}\ll R_{1},\ r_{n}\ll R_{n},\ \dots,\ r_{N}\ll R_{N}$, t-HOSVD assures a higher compression while still capturing the essential structure of the original tensor. However, the choice of parameters $(r_{1},r_{2},\dots,r_{N})$ directly influences the balance between compression (memory savings) and accuracy (quality of the reconstruction of the original tensor).
Another crucial aspect to consider when algorithms are used for compression is their computational efficiency in terms of FLOPs (floating point operations per second). In the HOSVD algorithm, SVD is used to decompose N flattened matrices $\mathbf{X}_{(n)}$, with $I_{n}\ll I$ $(I=\prod_{i\neq n}I_{i})$. Thus, the computational cost for these steps is given by the sum of N individual SVDs:
@@
In the context of data compression, assuming fixed accuracy for images or videos, the goal is to maximize compression or save memory space. Similarly, we introduce the tenSVD algorithm; however, we first want to focus on computational cost and its impact on energy consumption and processing time.
@@ 3.1 Computational Cost
From Section 2.3, it is easy to verify that the left singular vector of each $\mathbf{X}_{(n)}$ has $\mathcal{O}(I_{n}I^{2}+I^{3})$ FLOPs. On the other hand, it is clear that calculating the left singular vectors of $\mathbf{X}_{(n)}\mathbf{X}_{(n)}^{t}$ requires less computational overhead.
In fact, we need $\mathcal{O}(II_{n}^{2})$ FLOPs for multiple $\mathbf{X}_{(n)}$ with its transpose, and $\mathcal{O}(I_{n}^{3})$ for the SVD of $\mathbf{X}_{(n)}\mathbf{X}_{(n)}^{t}$. Thus, the total cost to obtain the left singular vectors is $\mathcal{O}(II_{n}^{2}+I_{n}^{3})$.
Following this approach, the FLOPs required by HOSVD are given by
$\sum_{n=1}^{N}\mathcal{O}(I_{n}I^{2}+I_{n}^{3}).$ (5)
Since $I_{n}\ll I$, $\mathcal{O}(I^{3})$ is greater than $\mathcal{O}(I_{n}^{3})$.
Based on this consideration, the primary concept of the new approach involves rearranging all elements of an Nth-order tensor $\mathcal{X}$ into an Mth-order tensor $\mathcal{Z}$ with $(M\ll N)$. Let $(I_{1},I_{2},\dots,I_{N})$ and $(J_{1},J_{2},\dots,J_{M})$ represent the dimensionality of $\mathcal{X}$ and $\mathcal{Z}$, respectively. In this transformation, we have $I_{n}\geq J_{m}$ for all m and n, ensuring that the total number of elements remains conserved through rearrangement, such that:
@@
Let $J=\prod_{j\neq m}J_{j}$. The computational cost of M matrices $\mathbf{Z}_{(m)}\mathbf{Z}_{(m)}^{t}$ is given by
$\sum_{m=1}^{M}\mathcal{O}(J_{m}J^{2}+J_{m}^{3}).$ (6)
It is easy to observe that $J_{m}J^{2}$ and $I_{n}I^{2}$ are the dominant terms in equations 6 and  5, which outweigh the cubic terms $J_{m}^{3}$ and $I_{n}^{3}$, respectively. Hence, the reordering of the elements across more modes reduces the computational cost from $\sum_{n=1}^{N}I_{n}I^{2}$ to $\sum_{m=1}^{M}J_{m}J^{2}$.

### v2

@@ Abstract
Tensors provide a robust framework for managing high-dimensional data. Consequently, tensor analysis has emerged as an active research area in various domains, including machine learning, signal processing, computer vision, graph analysis, and data mining. This study introduces an efficient image storage approach utilizing tensors, aiming to minimize memory to store, bandwidth to transmit and energy to processing. The proposed method organizes original data into a higher-order tensor and applies the Tucker model for compression. Implemented in R, this method is compared to a baseline algorithm. The evaluation focuses on efficient of the algorithm measured in term of computational time and the quality of information preserved, using both simulated and real datasets. A detailed analysis of the results is conducted, employing established quantitative metrics, with significant attention paid to sustainability in terms of energy consumption across algorithms.
journal:
@@ 1 Introduction
A large volume and complexity of image and video datasets are generated by modern devices every second. All this data necessitate continuous improvement of methods to reduce storage space and processing time, thereby facilitating more efficient extraction and transmission of information. It is well known that there are two major categories of methods for reduce memory space. The latter guarantees substantial reductions in memory cost while maintaining a high percentage of original information. Ongoing research endeavors persist in exploring and developing innovative methods to address the challenges associated with compression tasks.
@@
To this end, a diverse array of methodologies has been proposed across various disciplines, including astronomy, biology, chemistry, engineering, geology, and medical imaging. Several comprehensive surveys have been conducted to organize the different approaches employed in these domains, with [5] and [1] being notable recent surveys comparing several techniques for image and video compression.
@@
It is well established that image compression can be effectively achieved utilizing multivariate techniques such as Singular Value Decomposition (SVD) [11], [9]. While grayscale images are typically represented as two-dimensional matrices, color images are commonly structured in three dimensions (RGB channels). These multivariate methods can be applied to diminish the dimensionality of each color channel independently. However, such approaches become inefficient when dealing with videos or more complex signals that necessitate consideration of higher-dimensional data.
@@
Consequently, methodologies such as Tucker analysis [13] and broader tensor decompositions are increasingly leveraged for compression purposes. For an overview of tensor decomposition models, refer to [10]. These approaches are better equipped to manage the complexity inherent in high quality of images or videos and their convert to higher-dimensional data, thereby enhancing both compression efficiency and overall performance. A significant bottleneck for all compression methods is the need to achieve only the helpful information included in an high resolution images or videos. In this way a low storage footprint is need while maintaining an acceptable level of quality. The trade-off between reduced storage and the details and clarity of images or videos is consistently considered in compression contexts. Additionally, the time required for compression processes is a critical factor in signal transmission, as it directly influences transmission efficiency and bandwidth utilization. Rarely compression is studied as a fundamental tool for achieving in a more sustainable way images and videos. However, concerns regarding the energy consumption of data centers are rapidly increasing, and more of the huge amounts of energy that their consumption is for processing and archiving the million of images and videos that are collected every day. Thus, we must pay more attention when developing algorithms for image and video compression.
@@
Based on tensor decomposition, a novel algorithm is proposed for this aim, called tenSVD (tensor Singular Value Decomposition). It is designed to ensure high sustainability in terms of energy consumption but balancing the trade-off between storage and quality requirements.
@@
To ensure a fair comparison, processing time, particularly compression time with the same CPU, serves as an effective proxy for evaluation efficiency. Thus, tenSVD is compared to the established algorithm for Tucker analysis, Higher-Order Singular Value Decomposition (HOSVD) [3]. The strengths and weaknesses of tenSVD are analyzed in terms of several metrics, including Peak Signal-to-Noise Ratio (PSNR), Mean Squared Error (MSE), and Processing Time (TIME).
@@
The remainder of this article is structured as follows. Section 2 introduces tensor, including preliminary concept. Section 3 presents the main idea of approach proposed and relative algorithm. Case study with full comparative results are in Sections 4. Finally, in Section 5 are given conclusions and findings.
@@ 2 Tensor decomposition
Tucker model (Tucker3) is one of the most cited and applied methods for tensor decomposition. Originally proposed for three-way cases, it is referred to in several ways, including Three-Mode Principal Component Analysis (3MPCA), N-Mode Components Analysis (N-Mode PCA), and Higher-Order SVD (HOSVD) [13, 8, 6, 2]. Today, the Tucker model has been generalized to higher-order tensors and it is quickly increasing its application for imaging analysis [bip2021tensors]. For clarity, some preliminary concepts and notations will be recalled in Section 2.1 and will be used throughout the paper.
@@ 2.1 Preliminary concepts and notations
Objects with more than two indices are referred to as multi-way arrays, whereby the number of indices is termed the order. Nowadays, there is a trend to call any multi-way array as a tensor, and we keep this tradition to identify the linear operators with their multidimensional objects [12].
A real-valued tensor $\mathcal{X}\in\mathbb{R}$ of order $N\in\mathbb{N}$ is denoted by indices $(I_{1}\times\dots\times I_{N})$, and its entries by $x_{i_{1}\dots i_{N}}$. When only a fixed subset of indices is considered, this part of tensor is termed a subtensor. The fibers are defined by fixing every index except one, i.e. $\mathbf{x}(:,i_{2},i_{3},\dots,i_{N})$, and slices obtained by fixing all but two indices, i.e. $\mathbf{X}(:,:,i_{3},\dots,i_{N})$.
Tensors can be reformatting in several ways, and a particular case of manipulation is called n-mode unfolding or matrix unfolding. Denoted by $\mathbf{X}_{(n)}$, the n-mode unfolding of a tensor $\mathcal{X}$ arranges the n-mode fibers as columns in the resulting matrix: $\mathbf{X}_{(n)}\ (I_{n}\times\prod_{i\neq n}I_{i})$.
According to the Kolda and Bader [7], tensor products involved are significantly more complex compared to matrices.
Let be $\mathcal{X}(I_{1}\times\dots\times I_{n}\times\dots\times I_{N})$ and $\mathbf{U}(H\times I_{n})$,
@@
with $\mathcal{Y}(I_{1}\times\dots\times H\times\dots\times I_{N})$, and where $\times_{n}$ is the n-mode product of a tensor.
@@
The inner product of two same-sized tensors $\mathcal{X}$, $\mathcal{Y}$ is the sum of the products of their corresponding entries:
$\left\langle\mathcal{X},\ \mathcal{Y}\right\rangle=\sum_{i_{1}=1}^{I_{1}}\sum_{i_{2}=1}^{I_{2}}\dots\sum_{i_{N}=1}^{I_{N}}x_{i_{1}i_{2}\dots i_{N}}y_{i_{1}i_{2}\dots i_{N}}.$
@@ 2.2 Tucker model
In accord to the notation given in Section 2.1, the Tucker model for a Nth order tensor can be written as
@@
where each $\mathbf{U}^{(n)}(I_{n}\times R_{n})$ is an orthonormal component matrix ($\mathbf{U}^{(n)t}\mathbf{U}^{(n)}=\mathbf{I}$); and $\mathcal{G}$ is the core tensor with dimension $(R_{1}\times R_{2}\times\dots\times R_{N})$.
Assuming $N=2$ with dimension $(I_{1}\times I_{2})$, and $R_{1}$ and $R_{2}$ the dimension of vector space spanned by the first and second modes respectively, Tucker decomposition can be written as
@@
where $\mathbf{U}^{(1)}(I_{1}\times R_{1})$ and $\mathbf{U}^{(2)}(I_{2}\times R_{2})$ are orthonormal component matrices; and $\mathcal{G}$ is the core tensor with dimension $(R_{1}\times R_{2})$.
It is possible to show that $\mathbf{U}^{(1)}$ and $\mathbf{U}^{(2)}$ are the left and right singular vectors of $\mathcal{X}$, while the core tensor is a matrix with singular values on the first $r$ diagonal elements ($r=min(R_{1},R_{2})$) and 0s otherwise.
It is important underline that the core tensor for $N>2$ is a full tensor, where their elements can be negative too and their magnitude indicate the strength of interaction between components represented by the corresponding indices in the component matrices.
@@ 2.3 Tucker parameters estimation
The Tucker decomposition of a 2nd order tensor is equivalent to a SVD. Thus, it is not a surprise that HOSVD is the most popular algorithm for estimate the parameters of a Tucker model in case of $N>2$.
Introduced by De Lathauwer et al. [2], which generalizes the three-ways case presented by Tucker [13], HOSVD finds the exact Tucker decomposition of rank-$(R_{1},R_{2},\dots,R_{N})$, and subsequently provides the core tensor $\mathcal{G}$. In essence, HOSVD initially computes the left singular vectors of $\mathbf{X}_{(n)}$ for $n=1\dots N$. Finally, the transposes of these component matrices are multiplied with the original tensor $\mathcal{X}$ to obtain the core tensor.
When exact parameters are not required, as well as in case of compression, only the first $r_{n}$ columns of $\mathbf{U}^{(n)}$ are considerated, we need to solve the objective function
@@
where $\mathbf{U}^{(n)}(I_{n}\times r_{n})$ is orthonormal; and $\mathcal{G}$ has dimension $(r_{1}\times r_{2}\times\dots\times r_{N})$.
The details of the optimization problem are readily available [3, 7], when the version of HOSVD to estimate the parameters is called truncated HOSVD (t-HOSVD). It is briefly described below highlighting its simplicity and limited number of steps.
@@
Based on lossless compression strategy, HOSVD can reduce significantly the memory cost of the original tensor. Its effectiveness is determined by $\textrm{rank}-(R_{1},R_{2},\dots,R_{N})$ compared the dimensions of modes $(I_{1},I_{2},\dots,I_{N})$. In fact, original memory cost of $\mathcal{X}$ is $\prod_{n=1}^{N}I_{n}$, while the memory cost of HOSVD is given by a sum of smaller matrices and a core tensor, i.e $\sum_{n=1}^{N}I_{n}\times R_{n}$ elements of component matrices and $\prod_{n=1}^{N}R_{n}$ for core tensor.
Differently, by choose $r_{1}\ll R_{1}\dots r_{n}\ll R_{n}\dots r_{N}\ll R_{N}$, t-HOSVD assures an higher compression while still capturing the essential structure of the original tensor. However, the choice of parameters $(r_{1},\ r_{2},\dots,\ r_{N})$ directly influences the balance between compression (memory savings) and accuracy (quality of the reconstruction of the original tensor).
Another crucial aspect to consider when algorithms are used for compression is their computational efficiency in terms of FLOPs (floatting points operation per seconds). In the HOSVD algorithm, SVD is used to decompose $N$ flattened matrices $\mathbf{X}_{(n)}$, with $I_{n}\ll I$ ($I=\prod_{i\neq n}I_{i}$), thus computational cost for these steps is given by the sum of $N$ single SVDs:
@@
In data compression context, fixed the accuracy of image or video, the goal is maximize the compression or saving memory space. In same way we introduce tenSVD algorithm, but we want before pay attention on computational cost and its impact on energy and time consuming.
@@ 3.1 Computational cost
From Sections 2.3, it is easy verify that left singular vector of each $\mathbf{X}_{(n)}$ has $\mathcal{O}(I_{n}I^{2}+I^{3})$ FLOPs. On the other side, it is clear that calculating the left singular vectors of $\mathbf{X}_{(n)}\mathbf{X}_{(n)}^{t}$ requires less computational overhead. In fact, we need $\mathcal{O}(II^{2}_{n})$ FLOPs for multiple $\mathbf{X}_{(n)}$ with its transpose, and $\mathcal{O}(I_{n}^{3})$ for the SVD of $\mathbf{X}_{(n)}\mathbf{X}_{(n)}^{t}$. Thus, the total cost for obtain left singular vectors is $\mathcal{O}(II^{2}_{n}+I_{n}^{3})$.
Following this approach, the FLOPs required by HOSVD are
$\sum_{n=1}^{N}\mathcal{O}(I_{n}I^{2}+I^{3}_{n}).$ (5)
Due to $I_{n}\ll I$, $\mathcal{O}(I^{3})$ is higher then $\mathcal{O}(I^{3}_{n})$.
Based on this consideration, the primary concept of the new approach involves rearranging all elements of an Nth-order tensor $\mathcal{X}$, into an Mth-order tensor $\mathcal{Z}$ with $(M\ll N)$. Let be $(I_{1},I_{2},\dots,I_{N})$ and $(J_{1},J_{2},\dots,J_{M})$ the dimensionality of $\mathcal{X}$ and $\mathcal{Z}$ respectively. In this transformation, we have that $(I_{n}\geq J_{m})$ for all $m$ and $n$, and ensuring that the total number of elements remains conserved through rearrangement, such that:
@@
Let be $J=\prod_{j\neq m}J_{j}$, the computational cost of $M$ matrices $\mathbf{Z}_{(m)}\mathbf{Z}_{(m)}^{t}$ is
$\sum_{m=1}^{M}\mathcal{O}(J_{m}J^{2}+J^{3}_{m}).$ (6)
It is easy to observe that $J_{m}J^{2}$ and $I_{n}I^{2}$ are the dominant terms in equation 6 and  5, which outweighs the cubic term $J_{m}^{3}$ and $I_{n}^{3}$ respectively. Hence, the reordering of the elements across more modes, reducing the computational cost from $\sum_{n=1}^{N}I_{n}I^{2}$ to $\sum_{m=1}^{M}J_{m}J^{2}$.